# TopoGEN: topology-driven microstructure generation for in silico modeling of fiber network mechanics

### v1

@@ 4.2 Impact of microstructure on stress-strain behavior
We propose a novel computational approach to systematically analyze how experimentally measured microstructural features affect the nonlinear mechanical properties of the network. To achieve this, we subjected a library of topologically rich and heterogeneous RVEs to shear tests. In constructing the library, we incorporate the experimental evaluations reported by Jansen et al. (2018), wherein temperature variations induce a change in average valency by altering the number of cross-links and branches. Beyond influencing valency, temperature variations also induce morphological changes in the fibers, affecting their average length and diameter. We account for all such microstructural variations and summarize our findings in Table 1. Our results demonstrate that each microstructural parameter exerts a distinct influence on the differential modulus. The network-related parameters exhibit the most pronounced influence on the mechanical response. Specifically, we observe a homogeneous valency effect across the strain range tested, while the concentration effect is particularly prominent in the low-strain regime, consistent with previous findings (Sharma et al., 2016). Regarding fibril effects, the morphology of the fibrils causes only minor overall variations: in the medium-strain regime, a larger fibril radius makes the network stiffer, while a longer fibril softens it. In contrast, the intrinsic stiffness of the fibrils plays a significant role in modulating the variation of the differential modulus, especially in the low-strain regime. These results are important because they allow us to unravel the impact of various microstructural features that are influenced by environmental features such as temperature, concentration, and polymerization, highlighted by Andriotis et al. (2023) to have a crucial impact on the behavior of the collagen network. Indeed, while previous studies have focused on the extraction of network parameters from gels polymerized at different concentrations (Lindström et al., 2010, 2013), experiments show that concentration alone cannot fully characterize the network structure (i.e., mean valency) and the mechanisms of fibril formation (i.e., fibril length, radius, and stiffness). Our work presents a mechanistic approach for modeling discrete fiber networks that explicitly incorporates microstructural variability. This method is particularly valuable for informing experimental investigations, where the challenge lies in deciphering how individual microstructural features influence the overall mechanical response. This complexity arises from the interplay of multiple multiscale factors, which are difficult to disentangle using typical laboratory protocols. Moreover, because microstructural effects can evolve during rheological loading, typically manifesting as a nonlinear increase in global network stiffness, any comprehensive analysis must account for strain-dependent behavior. Our models thus offer a powerful tool to support experimental assessments of these intricate interactions. In our particular case, we focused on the influence of concentration and temperature, as reported by Jansen et al. (2018), but in the future, we could think of expanding the library by looking at the combined effect of these two parameters or the effect of new environmental parameters, such as pH. For assessing the topology and microstructure effect on differential shear modulus across different strain ranges we assume independence among all the structural parameters. We must underline that this assumption introduces limitations because we disregard the potential existence of confounding factors. Future studies need to systematically isolate confounding factors when analyzing the nonlinear elastic response of collagen or other types of biopolymer networks. Another assumption in our work is that concentration remains constant during polymerization. Advanced imaging could test whether temperature-induced affects fiber volume fraction and provide a clearer understanding of how concentration changes during polymerization. In this work, we focus exclusively on elastic effects, assuming that fiber connections remain permanent with no cross-link breakage. To further enhance the microstructural analysis, future studies should consider time-dependent effects, including the irreversibilities introduced by cross-link breakage.
@@ 4.3 Impact of microstructure on non-affinity
To simulate biopolymer networks polymerized under varying conditions, we optimize random networks with data derived from high-resolution images. While previous studies (Lindström et al., 2010, 2013; Nan et al., 2018; Eichinger et al., 2021a; Mahutga et al., 2023; Leng et al., 2021) treated networks with multiple fibers intersecting at a single node, our model takes a different approach. We exclusively consider fiber-fiber interactions as permanent cross-links or branches, deliberately avoiding higher-order connections that could stabilize the network locally. This choice aligns with the concept of non-affinity that has been increasingly highlighted in recent experimental (Cavinato et al., 2020) and computational (Mahutga et al., 2023; Chen et al., 2023; Parvez and Picu, 2023) works as a crucial factor in understanding local alterations in the extracellular space. Non-affinity arises from the loose interconnections (D’Amore et al., 2010; Burla et al., 2020) between fibers that fall below the Maxwell isostatic threshold (Maxwell, 1864). According to the Maxwell criterion, a three-dimensional central force network is stable only if the average number of fibers per node exceeds six. In isotropic networks, where nodes are branches or cross-links, each branch corresponds to three crossing fibers, and each cross-link corresponds to four. This configuration results in an average number of fibers per node that falls below the Maxwell threshold. In such networks, if the fibers behaved solely as axial load-bearing elements (e.g., springs), the structure would be prone to collapse. However, the structural integrity of networks of stiff filaments - such as collagen networks - is maintained either through intrinsic stabilizing effects of the fibers (e.g., bending rigidity) or extrinsic factors related to their surrounding environment (e.g., loading conditions). We reflect this principle in our computational model, where stability is ensured by (1) athermal collagen fibrils with a bending-to-stretching stiffness ratio on the order of $10^{-3}$, as demonstrated in Licup et al. (2015), and (2) boundary-applied strain, which stabilizes the system by transitioning it from a non-affine to an affine configuration where fibers align with the loading direction and deform consistently with the macroscopic deformation gradient. We define $NS$ as a metric to quantify the level of non-affinity by measuring the deviation of cross-link displacements obtained from simulations relative to those predicted under a purely affine deformation assumption. The outcome of this score is consistent with previous studies (Broedersz et al., 2011; Sharma et al., 2016), which demonstrate that non-affinity decreases with increasing average valency and applied strain. Additionally, the inconsistency between microscopic and macroscopic affine displacements is most evident at the strain extremes in the low-strain regime for both tested valencies. In this regime, theoretical models predict a more uniform load distribution across the network. This observation confirms that biological networks adopt a highly heterogeneous load distribution strategy and align with previous works (Chandran and Barocas, 2005; Cavinato et al., 2020; Mahutga et al., 2023). This loading heterogeneity is most pronounced during the low-to-medium-strain regime, where most of the fibers rotate or bend while just a few actually stretch. With these findings, we highlight the importance of heterogeneous three-dimensional models in accurately capturing localized phenomena in the space surrounding cells and provide an in-depth insight into the mechanisms of gradual fiber recruitment.
@@ 4.4 Untangling the network puzzle
We examine the influence of the microstructure on the shear differential modulus of collagen networks polymerized at various temperatures. Overall, for all the tested conditions, our analysis highlights a typical three-phase response in discrete fiber networks under load. In the linear elastic phase, the differential modulus remains constant as strain increases, and geometric nonlinearity is negligible. This is followed by the nonlinear stiffening phase, where fibers reorient along the loading direction. A subset of fibers, initially oriented perpendicularly, undergoes significant bending. This geometrically induced nonlinearity happens when non-affine, bending-dominated deformation modes develop. Finally, the stress-strain curve becomes linear in the affine-dominated phase. In this region, stress paths form through fiber chains that connect the boundary faces and limit further structural changes. Our results from the valency effect tests show trends consistent with in vitro rheological tests (Jansen et al., 2018), demonstrating a two-decade increase in stiffness from the initial to the final strain. However, the valency effect alone does not fully account for the observed phenomena. Specifically, networks with higher valency exhibited stiffer responses at low strains and a less pronounced nonlinear effect at higher strains in vitro. Yet, in our computational models and those of Jansen et al. (2018), higher valency networks exhibit nonlinear stiffening patterns that align more closely with those observed in lower valency networks. This suggests that nonlinear elastic behaviors in networks polymerized at different temperatures cannot be fully explained by valency alone. Our results in Figure 5 highlight the impact that fibril-related parameters such as length, fibril radius, and stiffness have. Another study (Davoodi Kermani et al., 2021) explored the effect of varying morphological and graph descriptors on the linear elastic modulus of different networks. While providing valuable insights, it lacks direct applicability to biological networks subjected to varying environmental conditions and shear loads, resulting in highly nonlinear stress buildup in rheological experiments. In our work, we focus solely on experimentally relevant microstructural variations (as summarized in Table 6.1) and evaluate the differential modulus as a function of strain. We propose a novel approach to modeling biologically reconstituted networks that strategically incorporates the variations introduced by polymerization conditions during their reconstruction. In parallel, this approach can also address tissue-specific variations in microstructures observed in-vivo, where other environmental factors can lead to localized changes in the microstructure that result in different mechanical responses. In both cases, it is essential to consider additional structural and topological features to untangle the network puzzle fully. We should direct significant efforts toward extracting detailed network topologies from high-resolution images captured under varying polymerization conditions. Doing so can enrich our models with robust statistical structural variations and provide deeper insights into the structure-to-mechanics relationship across multiple scales.
@@ 5 Conclusion
Our study provides a detailed analysis of the effects of topology and microstructure on the nonlinear differential shear modulus of biological networks. Without loss of generality for other biological and non-biological fiber networks, we use our developed framework to replicate the subisostatic microstructural architectures typical of collagen networks, enabling a quantitative evaluation of localized non-affine phenomena within these complex three-dimensional environments. We achieve this condition by optimizing the average valency of in silico networks virtually replicating the temperature modification effect on the topology. Additionally, we expand the range of studied conditions by assessing other microstructural features beyond valency. Specifically, we demonstrate that fiber morphology, mechanics, and concentration influence network mechanics across different deformation regimes. For instance, fibril stiffness and concentration have a more significant impact on the low-deformation regime. In contrast, fibril length and diameter play a more significant role in the medium-deformation regime. By aligning our computational models with experimentally derived microstructural characteristics, we propose a generative pipeline that, while rooted in network theory, is foremost designed for experimental application. This pipeline focuses solely on structural parameters that can be experimentally controlled and will benefit from further investigations into how topology changes under varying environmental conditions and, consequently, how these changes affect mechanical responses.
@@ Acknowledgement
This work was supported by the Delft University of Technology Startup Grant awarded to MP and BF. The authors thank Prof. F.Gijsen (Delft University of Technology, the Netherlands) and Prof. G. Koenderink (Delft University of Technology, the Netherlands) for many useful discussions.

### v2

@@ 4.2 Impact of microstructure on stress-stretch behavior, a case study of collagen networks
To assess the robustness of our generative pipeline, we subject a library of topologically rich RVEs to shear tests. In constructing the library, we incorporate the experimental evaluations reported by [43], wherein temperature variations induce a change in average connectivity by altering the number of cross-links and branches. Beyond influencing connectivity, temperature variations also induce morphological changes in the fibers, affecting their average length and diameter. We account for all such microstructural variations and summarize our findings in Figure 6. Consistent with trends reported in the literature, our results confirm that each microstructural parameter has a distinct influence on the differential modulus [43, 51, 85]. Specifically, our results reveal a uniform influence of connectivity across the stretch range, in agreement with prior work [43] identifying local connectivity as a key determinant of collagen network elasticity. In contrast, collagen concentration predominantly affects the low-stretch regime. This sensitivity of the low-stretch mechanical response to collagen concentration is aligned with prior theoretical predictions: [51] showed that, in this regime, the elastic energy is dominated by bending modes, and the linear shear modulus scales quadratically with the fiber volume fraction. The nonlinear stiffening, however, becomes essentially insensitive to network concentration at higher deformations, a behavior that appears to be unique to collagen [51] and fibrin networks [41]. To verify that the effect of collagen concentration on the low-stretch mechanical response is independent of network connectivity, we examine the differential modulus across varying concentrations and connectivity levels, consistently observing the same concentration-driven trends. Changes in fibril radius and length do not produce large shifts in the differential shear modulus, largely due to the limited parameter range explored, deliberately constrained to match experimental observations. Within this range, the bending-to-stretching stiffness ratio remains relatively constant, limiting the impact on network mechanics across stretch regimes. Nevertheless, even modest slenderness variations can lead to measurable stiffness changes, as network mechanics in the bending-dominated low-stretch regime is highly sensitive to fiber length and radius, in line with prior modeling studies [8, 39, 85]. In contrast, the intrinsic elastic modulus of the fibrils plays a significant role in modulating the variation of the differential modulus, especially in the low-stretch regime. The uneven low-to-high stretch regime effect of the fibril elastic modulus on the differential modulus - consistently observed across different polymerization conditions and network topologies - emerges from the imposed tension-compression fibril modulus asymmetry. This is supported by our simulations, where removing this intrinsic fiber elastic modulus asymmetry eliminates the effect. These results are important because they allow us to unravel the impact of various microstructural features that are influenced by environmental features such as temperature, concentration, and polymerization, highlighted by [1] to have a crucial impact on the behavior of the collagen network. Indeed, while previous studies have focused on the extraction of network parameters from gels polymerized at different concentrations [53, 52], experiments show that concentration alone cannot fully characterize the network structure (i.e., mean connectivity) and the mechanisms of fibril formation (i.e., fibril length, radius, and elastic modulus). Our work presents a mechanistic approach for modeling discrete fiber networks that explicitly incorporates microstructural variability. This method is particularly valuable for informing experimental investigations, where the challenge lies in deciphering how individual microstructural features influence the overall mechanical response. This complexity arises from the interplay of multiple multiscale factors, which are difficult to disentangle using typical laboratory protocols. Moreover, because microstructural effects can evolve during rheological loading, typically manifesting as a nonlinear increase in global network stiffness, any comprehensive analysis must account for stretch-dependent behavior. Our models thus offer a powerful tool to support experimental assessments of these intricate interactions. In our case study, we focused on the influence of concentration and temperature, as reported by [43], but in the future, we could think of expanding the library by looking at the combined effect of these two parameters or the effect of new environmental parameters, such as pH. For assessing the topology and microstructure effect on differential shear modulus across different shear ranges, we assume independence among all the structural parameters. We must underline that this assumption introduces limitations because we disregard the potential existence of confounding factors. Future studies need to systematically isolate confounding factors when analyzing the nonlinear elastic response of collagen or other types of biopolymer networks. Another assumption in our work is that concentration remains constant during polymerization. Advanced imaging could test whether temperature-induced affects fiber volume fraction and provide a clearer understanding of how concentration changes during polymerization. In this work, we focus exclusively on elastic effects, assuming that fiber connections remain permanent with no cross-link breakage. To further enhance the microstructural analysis, future studies should consider time-dependent effects across varying polymerization conditions, including the irreversibilities introduced by cross-link breakage.
@@ 4.3 Impact of microstructure on non-affinity
To simulate biopolymer networks polymerized under varying conditions, we optimize random networks with data derived from high-resolution images. While previous studies [53, 52, 64, 24, 55, 49] treated networks with multiple fibers intersecting at a single node, our model takes a different approach. We exclusively consider fiber-fiber interactions as permanent cross-links or branches, deliberately avoiding higher-order connections that fail to replicate the loosely connected architecture characteristic of collagen networks. This choice aligns with the concept of non-affinity that has been increasingly highlighted in recent experimental [13] and computational [55, 15, 68] works as a crucial factor in understanding local alterations in the extracellular space. Non-affinity arises from the loose interconnections [23, 9] between fibers that, in the case of networks with only central-force interactions, fall below the Maxwell isostatic threshold ([56]). According to the Maxwell criterion, a three-dimensional central force network is stable only if the average number of fibers per node exceeds six. In isotropic networks, where nodes are branches or cross-links, each branch corresponds to three crossing fibers, and each cross-link corresponds to four. This configuration results in an average number of fibers per node that falls below the Maxwell threshold. In such networks, if the fibers behave solely as axial load-bearing elements such as springs, or if bending modes are allowed within the fiber but not transmitted at the nodes by using pin-jointed cross-links [76], the structure is prone to collapse. However, the structural integrity of networks of stiff filaments - such as collagen networks - is maintained either through intrinsic stabilizing effects of the fibers (e.g., bending rigidity) or extrinsic factors related to their surrounding environment (e.g., loading conditions). We reflect this principle in our computational model, where stability is ensured by (1) athermal collagen fibrils with a bending-to-stretching stiffness ratio on the order of $10^{-3}$, as demonstrated in [51], and (2) boundary-applied stretch, which stabilizes the system by transitioning it from a non-affine to an affine configuration where fibers align with the loading direction and deform consistently with the macroscopic deformation gradient. We define $NS$ as a metric to quantify the level of non-affinity by measuring the deviation of cross-link displacements obtained from simulations relative to those predicted under a purely affine deformation assumption. The outcome of this score is consistent with previous studies [7, 85], which demonstrate that non-affinity decreases with increasing average connectivity and applied stretch. Additionally, the inconsistency between microscopic and macroscopic affine displacements is most evident in the low-stretch regime for both tested topologies. In this regime, theoretical models predict a more uniform load distribution across the network. This observation confirms that biological networks adopt a highly heterogeneous load distribution strategy and align with previous works [14, 13, 55]. This loading heterogeneity is most pronounced during the low-to-medium-stretch regime, where most of the fibers rotate or bend while just a few actually stretch. With these findings, we highlight the importance of heterogeneous three-dimensional models in accurately capturing localized phenomena in the space surrounding cells and provide an in-depth insight into the mechanisms of gradual fiber recruitment. Other non-affinity measurements have been proposed in the literature to quantify this important phenomenon from a computational perspective [76, 86]. In line with these methods, we compare the model-predicted displacements to those expected from a homogeneous continuum-scale affine deformation map. Future research should focus on systematically extracting quantitative measures of non-affinity in vitro.
@@ 4.4 Untangling the network puzzle
We examine the influence of the microstructure on the shear differential modulus of collagen networks polymerized at various temperatures. Overall, for all the tested conditions, our analysis highlights a typical three-phase response in discrete fiber networks under load. In the linear elastic phase, the differential modulus remains constant as stretch increases, and geometric nonlinearity is negligible. This is followed by the nonlinear stiffening phase, where fibers reorient along the loading direction. A subset of fibers, initially oriented perpendicularly, undergoes significant bending. This geometrically induced nonlinearity happens when non-affine, bending-dominated deformation modes develop. Finally, the stress-stretch curve becomes linear in the affine-dominated phase. In this region, stress paths form through fiber chains that connect the boundary faces and limit further structural changes. These results are consistent with well-established findings in the literature [6, 51, 73]. Furthermore, our results from the connectivity effect tests show trends consistent with in vitro rheological tests [43], demonstrating a two-decade increase in stiffness from the initial to the final stretch. However, the connectivity effect alone does not fully account for the observed phenomena. Specifically, networks with higher connectivity exhibited stiffer responses at low stretches and a less pronounced nonlinear effect at higher stretches in vitro. Yet, in our computational models and those of [43], higher connectivity networks exhibit nonlinear stiffening patterns that align more closely with those observed in lower connectivity networks. This suggests that nonlinear elastic behaviors in networks polymerized at different temperatures cannot be fully explained by connectivity alone. Our results in Figure 5 highlight the impact that fibril-related parameters such as length, fibril radius, and stiffness have. Another study [19] explored the effect of varying morphological and graph descriptors on the linear elastic modulus of different networks. While providing valuable insights, it lacks direct analysis of biological networks under varying environmental conditions and shear loads throughout the highly nonlinear stress buildup observed in rheological experiments. In our work, we focus solely on experimentally relevant microstructural variations (as summarized in Table 6.1) and evaluate the differential modulus as a function of the shear. We propose a novel approach to modeling biologically reconstituted networks that strategically incorporates the variations introduced by polymerization conditions during their reconstruction. In parallel, this approach can also address tissue-specific variations in microstructures observed in-vivo, where other environmental factors can lead to localized changes in the microstructure that result in different mechanical responses. In both cases, it is essential to consider additional structural and topological features to untangle the network puzzle fully. We should direct significant efforts toward extracting detailed network topologies from high-resolution images captured under varying polymerization conditions. Doing so can enrich our models with robust statistical structural variations and provide deeper insights into the structure-to-mechanics relationship across multiple scales.
@@ 5 Conclusion
Our study provides a detailed analysis of the effects of topology and microstructure on the nonlinear differential shear modulus of biological networks. Without loss of generality for other biological and non-biological fiber networks, we use our developed framework to replicate the loosely interconnected microstructural architectures typical of collagen networks, enabling a quantitative evaluation of localized non-affine phenomena within these complex three-dimensional environments. We achieve this condition by optimizing the average connectivity of in silico networks virtually replicating the temperature modification effect on the topology. Additionally, we expand the range of studied conditions by assessing other microstructural features beyond connectivity. Specifically, we demonstrate that even within a limited range of conditions - defined by structural variations in fiber morphology, their asymmetric response to tension-compression, and their concentration - these features influence network mechanics across different deformation regimes. By aligning our computational models with experimentally derived microstructural characteristics, we propose a generative pipeline that, while rooted in network theory, is foremost designed for experimental application. This pipeline focuses solely on structural parameters that can be experimentally controlled and will benefit from further investigations into how topology changes under varying environmental conditions and, consequently, how these changes affect mechanical responses.
@@ Acknowledgement
The authors thank F. Gijsen and G. Koenderink for useful discussions. This work was supported by the Delft University of Technology Startup Grant awarded to MP and BF, and the NWO Veni Talent Award 20058 to MP.